Log in
wiki.ucar.edu
Spaces
Hit enter to search
Help
Online Help
Keyboard Shortcuts
Feed Builder
What’s new
About Confluence
Log in
CABL Educational Field Project - EOL
Edit space details
Pages
Blog
Blog
2015
July
June
May
April
March
Status update Tuesday
Qsoil.ehs NA's
Changed heights on flux variables from 3 to 5 meters
Status update Monday
Daily update Sunday
bao sonic heights
Quickie status
Data outage last 2 days
BAO site visit
Tsoil.x.0.6cm.ehs Drifting
Quickie status
killed dhcpd on flux (again!)
ehs motes rewired
EHS Vmote.soil Problem
Quickie status
TP01 processing
Copied raw data files to HPSS
ehs down Mar 11 am
Mote sensor IDs
bao EC150 restarted
ehs installed
EC150 sonic and IRGA on bao flux station out since Mar 5 pm
How to check and restart processing
USB flash drive on 250m
Status Update Thursday
BAO surface station DSM questionable port 4
bao ec150 up
Status update Monday
February
Browse pages
Configure
Space tools
View Page
A
t
tachments (0)
Page History
Page Information
View Source
Export to PDF
Hide Inline Comments
Blog
2015
March
16
killed dhcpd on flux (again!)
Attachments
There are currently no attachments on this page.
Overview
Content Tools
{"serverDuration": 85, "requestCorrelationId": "37c385b741e33016"}